OpenStudy (anonymous):

domain of the function 1/sqrt(15x + 12) pls

OpenStudy (tkhunny):

Domain: Generally think of All Real Numbers and then see if you have to throw out anything. Common "throw-outs" are these. What makes a denominator zero? Logarithms need positive arguments. Even roots need non-negative arguments. Various other things, but these will do for this problem. What say you? What shall we discard?

OpenStudy (anonymous):

is it [78,inf)?

OpenStudy (tkhunny):

No. Far too much has been discarded. Tell me how you managed that result?

OpenStudy (anonymous):

15x+12=0 15x=-12 x=-12/15

OpenStudy (anonymous):

is it (-12/15,inf) then?

OpenStudy (tkhunny):

That is correct, but it is not a logical conclusion from your work. You need 15x + 12 > 0 This leads to x > -12/15 and is logically translated to \(\left(-\dfrac{12}{15},\infty\right)\)

OpenStudy (anonymous):

thanks very much

OpenStudy (anonymous):

what about this\[\sqrt[3]{t-78}\] what is the domain in interval notation please